What is ΔH° rxn (in kilojoules) for the reaction below?
3 Fe 2O 3(s) + CO(g) ↔ 2 Fe 3O 4(s) + CO 2(g)

Fe 2O 3(s) ΔH f o = -824.2 kJ/mol
CO(g) ΔH f o = -110.5 kJ/mol.
Fe 3O 4(s) ΔH f o = -1118.4 kJ/mol.
CO 2(g) ΔH f o = -393.5 kJ/mol.
1
Expert's answer
2020-12-17T07:33:49-0500

Q150909

What is ΔH° rxn (in kilojoules) for the reaction below?


3 Fe2O3(s) + CO(g) ↔ 2 Fe3O4(s) + CO2(g)


Fe2O3(s) ΔHf0 = -824.2 kJ/mol

CO(g) ΔHf0 = -110.5 kJ/mol.

Fe3O4(s) ΔHf0 = = -1118.4 kJ/mol.

CO2(g) ΔHf0 = = -393.5 kJ/mol.


Solution:


The given reaction is already balanced for all the element, so there is no need for balancing the

reaction.


The heat of reaction / Enthalpy of reaction: Heat of reaction is equal to the difference between the sum of enthalpies of products and the sum of enthalpies of reactants.


ΔH° rxn = Sum of enthalpies of products - Sum of enthalpies of reactants


ΔH° rxn = ΣaH0products - ΣbH0 reactants


a and b are the coefficients of the reactants and products in the balanced equation.


ΔH° rxn = ( 2 * ΔHF of Fe3O4 + 1 * Δ HF of CO2 )– ( 3 * ΔHF of Fe2O3 + 1 * Δ HF of CO )


ΔH° rxn = ( 2 * – 1118.4 kJ/mol + 1 * – 393.5 kJ/mol ) – ( 3 * – 824.2 kJ/mol + 1 * – 110.5 kJ/mol )


ΔH° rxn = ( – 2236.8 – 393.5 ) – ( – 2472.6 – 110.5)


ΔH° rxn = (– 2630.3)– (– 2583.1)


ΔH° rxn = -2630.3 + 2583.1 = -47.2 kJ ;


Hence ΔH° rxn (in kilojoules) for the given reaction is -47.2 kilojoules.